# A low-rank solver for conforming multipatch Isogeometric Analysis

### v3

@@ 4.3 Ring domain
Table 3 collects the number of iterations for degrees $p=2,3,4,5$ and number of elements per patch per parametric direction $n_{el}=2^{L}$ with $L=5,6,7,8$, as in the previous test. The number of iterations is independent of the meshsize $h$ and appears to be stable with respect to the polynomial degree $p$, although the case $p=2$ shows higher iteration counts compared to the higher degrees.
@@
$n_{el}$ $p=2$ $p=3$ $p=4$ $p=5$
32 59 42 39 40
64 63 44 41 42
128 71 50 46 46
256 77 60 58 55
Table 3: Number of iterations for the ring domain.
We report in Figure LABEL:fig:rank_ring the maximum of the multilinear ranks of the solution in all the directions and for all the subdomains. Interestingly, for degrees 2 and 4 this maximum rank appears to grow in the left part of of the plot (corresponding to small values of $n_{el}$), but it then decreases as $n_{el}$ grows and on the finest discretization level in both cases it reaches a value between 60 and 70, which is also (roughly) the maximum rank observed for degree 3.
Nevertheless, as represented in Figure LABEL:fig:mem_ring, the memory storage for the solution is hugely reduced with respect to the full solution and reaches values around $1\%$ for the highest discretization level.
Figure 7: Results for the ring domain.
We perform an additional test using the same geometry, boundary conditions and forcing term, considering non-constant Lamé parameters. In particular, in (4.1) we set $\nu=0.3$ and $E=1+\frac{1+\sin(\theta(x,y))}{2}$, where $\theta(x,y)$ is the argument of the complex number $x+iy$. The numerical solution is represented in Figure 8.
Figure 8: Numerical solution for the ring domain (variable Lamé parameters).
The numbers of TPCG iterations are reported in Table 4, while the maxima of the multilinear ranks of the solution and memory requirements are shown in Figure 9, for different different degrees and numbers of elements. The results are similar to the case with constant parameters.
 Iteration number
$n_{el}$ $p=2$ $p=3$ $p=4$ $p=5$
32 71 46 43 45
64 77 48 45 47
128 84 53 51 51
256 96 70 65 65
Table 4: Number of iterations for the ring domain (variable Lamé parameters).
Figure 9: Results for the ring domain (variable Lamé parameters).
@@ 4.4 Igloo-shaped domain
[…]
Figure 10: Initial geometry and numerical solution for the igloo-shaped domain.
@@
for $i,j=1,\dots,4$ s.t. $\Theta^{(i)}\cap\Theta^{(j)}$ has zero measure. Indeed, in Table 5 we reported the number of iterations for degrees $p=2,3,4$ and a number of elements per parametric direction $n_{el}=2^{L}$ with $L=4,5,6$. Note that the maximum values of $L$ and $p$ is slightly smaller than in the two previous tests. This is due to the higher multilinear ranks of this geometry. Nevertheless, note that the number of degrees of freedom for the finest discretizations level is higher than 7 millions.
@@
The maximum of the multilinear ranks of the solution in all the directions and for all the subdomains are represented in Figure LABEL:fig:rank_igloo.
The gain in terms of storage may seem less significant than in the previous cases, but it is in fact very similar if the comparison is done for the same discretization level. For this problem, on the finest discretization level, the memory storage of the low-rank solution is less than 35% that of the full solution.
Figure 11: Results for the igloo domain.
@@ 5 Conclusions
In this work, we have considered the isogeometric discretization of a compressible linear elasticity problem. We have extended the low-rank solution method presented [21], which is based on a low-rank Tucker representation of vectors and matrices, to conforming multipatch discretizations. This solver exploits the domain decomposition paradigm, where the subdomains are built from union of pairs of adjoining patches. Each block of the system matrix and each block of the right hand side vector is approximated by a Tucker matrix and a Tucker vector, respectively. The resulting linear system is solved by a truncated preconditioned conjugate gradient method. The designed preconditioner has a block-diagonal structure where each block is a Tucker matrix and its application exploits the fast diagonalization method and the fast Fourier transform.
We performed numerical tests assessing a low memory storage, almost independent of the degree $p$, and the number of iterations independentof $p$ and the mesh-size.
The problems considered are simple academic benchmarks, and we emphasize that the effectiveness of these techniques is greater when the rank of the considered approximation of the problem’s components (the geometry, the operator coefficients, the forcing term, and therefore the solution) is lower. The promising results obtained suggest further exploring, in future work, the potential of the proposed method in realistic applications. The method relies on the linearity of the problem, however, nonlinear problems could be addressed by combining outer Newton-type solvers with inner solvers of the proposed type for the linearized step.
Our tests also demonstrate that our procedure achieves the optimal accuracy expected by the isogeometric scheme. However, this paper does not delve into the theoretical error study. Beyond the convergence theory of truncated iterative solvers, which is a significant open research area, there is also the fundamental issue of understanding how the low-rank approximation of the problem and the subsequent approximations introduced by the solver impact the accuracy of the numerical solution. This study is non-trivial, given that low-rank truncations are computed measuring the approximation w.r.t. the euclidean norms, which is not the natural norm for the well-posedness of the (discretized) differential equation. The topic warrants further investigation and will be the focus of future research.
@@ Acknowledgments
he authors are members of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INDAM), the first author was partially supported by INDAM-GNCS Project “Sviluppo di metodi numerici innovativi ed efficienti per la risoluzione di PDE”, while the third author was partially supported by INDAM-GNCS Project “Nodi, ascisse e punti: scegliere e usare in maniera efficiente”. The second author acknowledges support from PNRR-M4C2-I1.4-NC-HPC-Spoke6. The third author acknowledges the contribution of the Italian Ministry of University and Research (MUR) through the PRIN project COSMIC (No. 2022A79M75). The second and the third author acknowledge support from the PRIN 2022 PNRR project NOTES (No. P2022NC97R).

### v4

@@ 5.4 Thick ring domain
The relative errors in 2-norm $\|\cdot\|_{2}$ of the approximation of the linear system matrix $\mathbf{A}$ by $\widetilde{\mathbf{A}}$ for $p=3,4,5$ and $n_{el}=8,16$ are reported in Table 4. These computations require the assembly of both matrices, which forced us to consider a limited number of elements per parametric direction, because of memory constraints. The resulting error is of the order of the Chebfun3F tolerance, which we recall is set equal to $10^{-7}$ for all the considered cases.
 $\|\mathbf{A}-\widetilde{\mathbf{A}}\|_{2}/\|\mathbf{A}\|_{2}$
$n_{el}$ $p=3$ $p=4$ $p=5$
8 $4.9\cdot 10^{-8}$ $3.6\cdot 10^{-8}$ $3.2\cdot 10^{-8}$
16 $9.1\cdot 10^{-8}$ $8.3\cdot 10^{-8}$ $7.6\cdot 10^{-8}$
Table 4: Relative errors of the approximation of $\mathbf{A}$ by $\widetilde{\mathbf{A}}$ in the thick ring domain.
Table 5 collects the number of iterations for degrees $p=3,4,5$ and number of elements per patch per parametric direction $n_{el}=2^{L}$ with $L=5,6,7,8$, as in the previous test. The number of iterations is independent of the meshsize $h$ and appears to be stable with respect to the polynomial degree $p$.
@@
$n_{el}$ $p=3$ $p=4$ $p=5$
32 42 39 40
64 44 41 42
128 50 46 46
256 60 58 55
Table 5: Number of iterations for the thick ring domain.
We report in Figure LABEL:fig:rank_ring the maximum of the multilinear ranks of the solution in all the directions and for all the subdomains. As represented in Figure LABEL:fig:mem_ring, the memory storage for the solution is hugely reduced with respect to the full solution and reaches values around $1\%$ for the highest discretization level.
Figure 10: Results for the thick ring domain.
@@ 5.5 Igloo-shaped domain
In this test we consider an igloo-shaped domain, represented in Figure 11. As in the previous test case, we have four patches and four subdomains, represented by the union of green and yellow patches, yellow and blue patches, blue and red patches and red and green patches. We impose homogeneous Dirichlet boundary conditions on $\{z=0\}$ and the internal boundary, while homogeneous Neumann boundary conditions are imposed elsewhere.
Figure 11: Igloo-shaped domain.
@@
for $i,j=1,\dots,4$ s.t. $\Theta^{(i)}\cap\Theta^{(j)}$ has zero measure. Also in this testcase, we report in Table 6 the relative errors in 2-norm $\|\cdot\|_{2}$ of the approximation of the linear system matrix $\mathbf{A}$ by $\widetilde{\mathbf{A}}$ for $p=3,4$ and $n_{el}=8,16$. Even in this case, the approximation is of the order of the Chebfun tolerance, which is set equal to $10^{-7}$ for all the cases considered.
 $\|\mathbf{A}-\widetilde{\mathbf{A}}\|_{2}/\|\mathbf{A}\|_{2}$
$n_{el}$ $p=3$ $p=4$
8 $3.7\cdot 10^{-8}$ $3.3\cdot 10^{-8}$
16 $7.0\cdot 10^{-8}$ $6.7\cdot 10^{-8}$
Table 6: Relative errors of the approximation of $\mathbf{A}$ by $\widetilde{\mathbf{A}}$ in the igloo-shaped domain.
In Table 7 we reported the number of iterations for degrees $p=3,4$ and a number of elements per parametric direction $n_{el}=2^{L}$ with $L=3,4,5,6$. Note that the maximum values of $L$ and $p$ are slightly smaller than in the two previous tests. This is due to the higher multilinear ranks of this geometry. Nevertheless, note that the number of degrees of freedom for the finest discretizations level is higher than 7 millions.
@@
The maximum of the multilinear ranks of the solution in all the directions and for all the subdomains are represented in Figure LABEL:fig:rank_igloo. The gain in terms of storage may seem less significant than in the previous cases, but it is in fact very similar if the comparison is done for the same discretization level. For this problem, on the finest discretization level, the memory storage of the low-rank solution is less than 35% that of the full solution.
Figure 12: Results for the igloo-shaped domain.
@@ 6 Conclusions
In this work, we have considered the isogeometric discretization of a compressible linear elasticity problem. We have extended the low-rank solution method presented [34], which is based on a low-rank Tucker representation of vectors and matrices, to conforming multipatch discretizations. This solver exploits the domain decomposition paradigm, where the subdomains are built from union of adjoining patches. Each block of the system matrix and each block of the right hand side vector is approximated by a Tucker matrix and a Tucker vector, respectively. In particular, we showed an upper bound for the relative error of the matrix stemming from this approximation. The resulting singular linear system is solved by a truncated preconditioned conjugate gradient method. The designed preconditioner has a block-diagonal structure where each block is a Tucker matrix and its application exploits the fast diagonalization method and the fast Fourier transform.
We performed numerical tests assessing a low memory storage, almost independent of the degree $p$, and the number of iterations independent of $p$ and the mesh-size. The problems considered are simple academic benchmarks, and we emphasize that the effectiveness of these techniques is greater when the rank of the considered approximation of the problem’s components (the geometry, the operator coefficients, the forcing term, and therefore the solution) is lower. The promising results obtained suggest further exploring, in future work, the potential of the proposed method in realistic applications. The method relies on the linearity of the problem, however, nonlinear problems could be addressed by combining outer Newton-type solvers with inner solvers of the proposed type for the linearized step.
Our tests also demonstrate that our procedure achieves the optimal accuracy expected by the isogeometric scheme. However, this paper does not delve into the theoretical error study. To demonstrate an error estimate for the solution for the error stemming from the use of a low-rank approximation of the data and the differential operator of the problem, as well as from the low-rank approximation of the variables constructed by the iterative solver is of primary importance to provide the mathematical justification for the use of low-rank techniques applied to solutions of partial differential equations in general, not just in the multipatch case. A complete error theory of this kind is not yet available and represents a central aspect of the theoretical development of these techniques. The topic warrants further investigation and will be the focus of future researches.
@@ Acknowledgments
The authors are members of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INDAM), the first author was partially supported by INDAM-GNCS Project “Sviluppo di metodi numerici innovativi ed efficienti per la risoluzione di PDE”, while the third author was partially supported by INDAM-GNCS Project “Nodi, ascisse e punti: scegliere e usare in maniera efficiente”. The second author acknowledges support from PNRR-M4C2-I1.4-NC-HPC-Spoke6. The third author acknowledges the contribution of the Italian Ministry of University and Research (MUR) through the PRIN project COSMIC (No. 2022A79M75). The first author acknowledges support from the PRINN 2022 PNRR project HEXAGON (No. P20227CTY3). The second and the third author acknowledge support from the PRIN 2022 PNRR project NOTES (No. P2022NC97R).